OpenStudy (anonymous):

Factor each polynomial, Identify the quadratic factor if one exists. x^7+x^7/2+x^5

OpenStudy (saifoo.khan):

\[\frac{2x^7}{2+x^5}\]??????????

OpenStudy (anonymous):

wait..i suck at equation drawing

OpenStudy (saifoo.khan):

u suck at everything.

OpenStudy (anonymous):

\[x ^{7}+x ^{7/2}+x^5\]

OpenStudy (anonymous):

explain it too :p

OpenStudy (saifoo.khan):

\[\Huge x^\frac72 (x^\frac72+x^\frac32+1) \]

OpenStudy (anonymous):

so u took out an x and a 7/2 rite?

OpenStudy (anonymous):

\[3r+7\sqrt{r}=11\]

OpenStudy (saifoo.khan):

Yes i took 7/2

OpenStudy (anonymous):

kay..how do u put tht in quadratic form?

OpenStudy (saifoo.khan):

umm.. im not sure about this one.

OpenStudy (anonymous):

kay..i will find it online sumwhere thanks for the help